AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 188     Entry time: Mon Feb 7 08:09:50 2022
Author: OH, CB and PH 
Subject: CARME Server setup and initial tests 
09:09 Arrived at GSI for comissioning beamtime.
      CARME currently in a stopped state, with all FEEs powered off.
      Upper detector is biased with 80V and has a leakage current of 9.92uA.
      This is up on the previous recording by JM of 5.4uA recorded in elog 187.
      It is possible light is reaching the detector via the interaction chamber which shall be checked.

09:19 Cave checked:
      Water is running to the FEEs (OK for power on when required)
      There are 4 uncovered windows on the interaction chamber. Will see if we can get these covered.

14:29 Tested different bias voltages.
      Between 80V and 120V got a normal diode response with small increases in leakage current
      At 120V leakage current was stable at 9.92 for one minute.
      Checked waveforms on DAQ and looked at current again and it was up to 16uA.
      Turned off bias and let ramp down.
      Now when trying to ramp up detector in similar position to last week with a diode response between 0-30V and a rapid increase above that.
      Carlo spoke to Gleb who was beam tuning. Which around the time of our trip he opened the valves and put beam into the ring for a rotation.
      Gleb put in several faraday cups prior to the CARME chamber yet we still see the non-diode bias characteristics.
      He will enter in 30 minutes and we will check if anything looks out of place. Will also remove motor pin to give option of movement.

15:00 Gleb and Carlo enters the cave.
      Dipole and beam off.
      Pin removed.
      Diode response restored 9.4uA at 40V. Which is only slightly above where it was during the morning


      Before they got anywhere with starting again the current has gone up again..

16:49 Since around 15:30 the current has come down to its levels earlier today at 40V. We are preparing a script to monitor the current levels going forward.

17:00 Gleb stopped tuning the beam for the day.
      There is no beam in the ring.
      Magnets and everything else continue as if there was beam in the ring.

19:40 Attempted to raise bias to 50. Immediately started to trip
      Carlo cycled motors multiple times.
      Bias in out position dropped.
      To lower
      Have made it to 60V. Will leave there overnight with leakage current tracking running.

      No beam so no daq running.
ELOG V3.1.4-unknown